# Model Extraction Attacks Revisited

### v1

@@ 1. Introduction
The remarkable advances in machine learning (ML) technologies in recent years (He et al., 2016; Deng and Platt, 2014; Roy et al., 2018; Naumov et al., 2019; Feng et al., 2021; Lagler et al., 2013; Zhou et al., 2023) have spurred the expansion of Machine-Learning-as-a-Service (MLaaS), which meets the increasing need for ML capabilities among users who might not possess the requisite expertise or infrastructure. MLaaS platforms offer publicly accessible APIs, enabling users to interface with backend ML models and systems seamlessly. Users are often charged on a per-query basis, making advanced ML capabilities both accessible and affordable. Many IT giants (e.g., Google, Amazon, Microsoft, and Face) have unveiled their MLaaS platforms.
However, there exists an inherent conflict between the proprietary nature of ML models and the public accessibility of MLaaS APIs, leading to a range of security concerns. One of the most prominent concerns is the threat of model extraction (ME) attacks. As illustrated in Figure 1, such attacks construct a piracy model functionally equivalent or similar to the backend model of MLaaS, via querying the black-box APIs with carefully crafted inputs. Since its conceptualization in (Tramèr et al., 2016), a stream of work has been proposed to improve the efficacy of ME attacks (Tramèr et al., 2016; Orekondy et al., 2019a; Krishna et al., 2019; Chandrasekaran et al., 2020; Pal et al., 2019, 2020; Pengcheng et al., 2018; Shi et al., 2018; Juuti et al., 2019; Jagielski et al., 2020; Yu et al., 2020; Truong et al., 2021; Carlini et al., 2021; Papernot et al., 2016a, b, 2018, 2017).
@@
Previous Conclusion |
Refined Conclusion |
Consistency |
Identifying the architectures of victim models and using the same architectures in piracy models substantially boost the effectiveness of ME attacks (Chen et al., 2022b). |
The architectures of piracy models have a limited impact on the performance of ME attacks, while more advanced models may not lead to more effective attacks. |
Using more complex models tends to achieve better attack performance (Shi et al., 2017; Krishna et al., 2019; Aïvodji et al., 2020). |
The impact of model complexity varies with the concrete tasks (i.e., FER versus NLU) while pre-training is a more dominating factor. |
Semi-supervised learning substantially improves the query efficiency of ME attacks, especially when the query budget is low (Jagielski et al., 2020). |
Semi-supervised learning improves the query efficiency of ME attacks, but the margin of improvement is not as large as reported in local experiments (Jagielski et al., 2020); further, it has a negative impact on adversarial fidelity. |
Active learning substantially improves the query efficiency of ME attacks (Pal et al., 2020). |
Active learning yields only marginal improvements in query efficiency and can, in some cases, have a negative effect. |
Using adversarial examples improves both the query efficiency and the attack effectiveness (Papernot et al., 2017; Juuti et al., 2019; Pengcheng et al., 2018; Yu et al., 2020). |
Using adversarial examples has a limited or even negative impact on attack fidelity, but may improve adversarial fidelity. |
Perturbing output confidence scores effectively mitigates ME attacks (Orekondy et al., 2019b; Lee et al., 2019). |
Output quantization weakens ME attacks but is not sufficiently effective. |
– Leveraging MeBench, we conduct an empirical study on leading MLaaS platforms (i.e., Amazon, Microsoft, Face, and Google) in the tasks of facial emotion recognition (FER) and natural language understanding (NLU). We show that today’s real-world MLaaS APIs still exhibit significant vulnerability to ME attacks, while the characteristics of vulnerability vary greatly across different platforms and tasks.
@@ 3. Status quo of ME vulnerability
Q: How vulnerable are today’s MLaaS APIs to ME attacks?
Q: How do various factors influence such vulnerability?
Q: How does the vulnerability evolve over time?
Q: How effective are the defenses (if any) employed on the MLaaS platforms?
@@ 3.1. Experiment setting
APIs – We consider the APIs of 4 leading MLaaS service providers: Amazon, Google, Microsoft, and Face. By default, we assume the adversary has access to the complete query response (i.e., classification labels and confidence scores). Specifically, in the FER task, each query response contains the confidence scores of different expressions (e.g., “calm”). In the NLU task, each query response includes the confidence scores of different sentiments (e.g., “positive”). The details of each API are deferred to Table 14.
@@ 3.2.2. Comparison with prior work
MLaaS APIs – Among prior work, Cloudleak (Yu et al., 2020) conducts ME attacks against MLaaS APIs and measures attack fidelity with respect to different APIs and query budgets. Specifically, it reports 81% fidelity and 58% accuracy against the Face API on the KDEF dataset. Furthermore, it reports that the attacks against other platforms such as Microsoft and Clarifai, as well as different tasks including traffic sign recognition, flower, and NSFW, all attain over 82% fidelity and 70% accuracy. This clearly showcases the widespread nature of this vulnerability across various APIs and tasks.
In our experiments, on the same platform and dataset (i.e., KDEF and Face), we observe that the attack attains similar results (74% fidelity and 66% accuracy), while the performance gap with (Yu et al., 2020) may be attributed to the scratch model and the difference of testing data. In (Yu et al., 2020), the testing data is synthesized, while we randomly sample 20% of the KDEF dataset as the testing set. Further, similar to (Yu et al., 2020), the ME vulnerability varies considerably across different APIs. In particular, the attack attains over 80% fidelity against the Microsoft API across all the datasets, while the other APIs demonstrate more variations from one dataset to another.
@@ 3.3.1. Optimizers
We specifically consider 4 popular optimizers: SGD (Robbins and Monro, 1951), Adam (Kingma and Ba, 2017), AdamW (Loshchilov and Hutter, 2019), and Lion (EvoLved Sign Momentum) (Chen et al., 2023), which is a recently proposed optimizer. We configure the basic ME attack with varied optimizers and evaluate its effectiveness in the FER and NLU tasks.
@@
FER – We compare different optimizers using RAFDB on the Amazon and Microsoft APIs. Following (Chen et al., 2023), we set the hyper-parameters (e.g., learning rate and decay rate ) for different optimizers as follows: SGD with and , Adam/AdamW with and , and Lion with and . We have the following key observations.
Table 3 shows the attack fidelity at the end of 200 training epochs. Under the same query budget, Lion, Adam, and AdamW attain comparable attack fidelity. There is no distinct advantage between Lion and AdamW; however, it is evident that SGD lags behind the performance of other optimizers significantly. We further examine the training dynamics of different optimizers. Figure 4 summarizes how the attack fidelity varies with the number of training epochs. Despite the claimed superior convergence speed of Lion in supervised learning (Chen et al., 2023), we observe that Adam/AdamW actually converges much faster than Lion in the ME attack, while SGD converges fairly slowly.
NLU – We also set the hyper-parameters of optimizers following (Chen et al., 2023): SGD with and , Adam/AdamW with and , and Lion with and .
Table 3 summarizes the attack performance corresponding to different optimizers on the IMDB dataset. Lion generally outperforms the other optimizers on both Microsoft and Amazon APIs. Our observations corroborate prior findings (Chen et al., 2023) that Lion outperforms alternative optimizers in fine-tuning large language models.
@@ 3.3.2. Piracy models
Model family – For FER, we compare 5 popular architectures, GoogLeNet (Szegedy et al., 2014), DenseNet (Huang et al., 2018), ResNet (He et al., 2016), EfficientNet (Tan and Le, 2020), Alex -Net (Krizhevsky, 2014), using RAFDB on the Microsoft and Face APIs, with results shown in Figure 5. For NLU, we compare 2 popular LLMs, RoBERTa and XLNet, using IMDB on the Amazon and Microsoft APIs, with results summarized in Table 5. We have the following interesting findings.
@@
Model complexity – We further assess the impact of model complexity (within a single model family). In FER, we use ResNet as the piracy model and vary its width and depth, spanning across ResNet18, ResNet50, ResNet101, and Wide-ResNet50. As illustrated in Figure 6, the attack fidelity across different architectures differs by less than and on Amazon and Face, respectively. Intuitively, as the model complexity reaches a certain level, further increasing the model width or depth does not lead to a notable improvement in attack performance. Similarly, in NLU, as shown in Table 5, for both RoBERTa and XLNet, using large models yields marginal improvement over base models. It is worth pointing out that compared with the base models, the large models of RoBERTa and XLNet are pre-trained using more data ( additional data) (Yang et al., 2019; Liu et al., 2019).
@@
In FER, we compare the attacks using a randomly initialized ResNet50 model as the piracy model with that using ResNet50 models pre-trained on public datasets. We consider two pre-training datasets, ImageNet-1K and VggFace2 (Cao et al., 2018). As shown in Figure 7, pre-training marginally improves the attack performance, while the improvement margin varies with the APIs. For instance, compared with VGGFace2, ImageNet-1K leads to more improvement on Microsoft but less so on Face. In NLU, we compare randomly initialized XLNet and RoBERTa and that pre-trained on the BERT dataset (Devlin et al., 2018), with results shown in Table 6. Surprisingly, unlike the PER task, pre-training language models significantly improves the attack performance. For instance, the pre-training of RoBERTa boosts fidelity by over 25%.
@@ 3.3.3. Attack strategies
It is observed that against the Amazon API, using PGD or CW to enhance ME attacks leads to a considerable improvement in attack fidelity (around 5.61% on average) on FER+, while the effect on RAFDB is fairly mixed. Meanwhile, against Face, adversarial learning-enhanced ME attack seems not only ineffective but even under-performs the basic attack. However, across all the cases, adversarial learning leads to a consistent improvement in adversarial fidelity. Our findings indicate that clean queries (sampled from the reference dataset) and adversarial queries (generated by adversarial attacks) seem to contribute to ME attacks differently, respectively improving the fidelity and adversarial fidelity of piracy models.
@@ 3.5. Q4: Defenses
It is unclear what specific defense mechanisms are employed by the MLaaS platforms. However, when examining their APIs, Amazon, Microsoft, and Face will provide exact confidence values for each class. In contrast, Google’s FER categorizes its output using five distinct likelihood descriptors, (i.e., “very_unlikely”, “unlikely”, “possible”, “likely” and “very_likely”) offering a different approach to presenting its analysis. This approach seems to be an attempt to obscure the confidence levels of the model’s output. The central concept behind such quantization is to reduce the granularity of the confidence values returned by the API, thereby making it more challenging for potential attackers to discern intricate details about the model’s internal processes. We mimicked Google’s approach and assess the effectiveness of quantization by applying it to other APIs. Specifically, We achieve this by discretizing the confidence values into intervals and substituting the confidence value with the midpoint of the corresponding interval (e.g., is represented as ). We manually reduce the granularity of the returned confidence information to examine its impact on the performance of the Model Extraction (ME) attack.
@@
| Face | +0.430.08 | -0.760.18 | +0.310.12 | -0.980.08 |
Table 13 presents the analysis results. When applied to the Amazon API, quantization resulted in an average decrease of 1.68% in the attack fidelity. In contrast, the Face API exhibits more mixed results in mitigating ME attacks.
@@ A.2. Datasets
| Task |
Dataset |
Size |
Description |
| F E R |
KDEF |
562*762 |
The exaggerated facial expressions made by subjects |
RAFDB |
100*100 |
Great-diverse facial images downloaded from the Internet |
EXPW |
224*224 |
Wild human facial expressions captured at social events |
FER+ |
48*48 |
Black and white human face expressions |
| N L U |
IMDB |
234 words |
Binary sentiment classification dataset, highly polar movie reviews collected from the internet movie database (IMDB) |
Yelp |
140 words |
Information from user reviews of different restaurants on Yelp. Ratings can be from 1-5, not a polarized dataset |

### v2

@@ 1. Introduction
The remarkable advances in machine learning (ML) technologies in recent years (Zhou et al., 2023; Zhang et al., 2021; Chen et al., 2023b; Li et al., 2024; Liang et al., 2021; Qi et al., 2023; Xu et al., 2023) have spurred the expansion of Machine-Learning-as-a-Service (MLaaS) (Li et al., 2022; Shi et al., 2018), which meets the increasing need for ML capabilities among users who might not possess the requisite expertise or infrastructure. MLaaS platforms offer publicly accessible APIs (Li et al., 2022; Feng et al., 2021), enabling users to interface with backend ML models and systems seamlessly. Users are often charged on a per-query basis, making advanced ML capabilities both accessible and affordable. Many IT giants (e.g., Google, Amazon, Microsoft, and Face++) have unveiled their MLaaS platforms.
However, there exists an inherent conflict between the proprietary nature of ML models and the public accessibility of MLaaS APIs, leading to a range of security concerns (Liu et al., 2023a, 2022; Li et al., 2019, 2023, 2021; Liu et al., 2024, 2023b; Lyu et al., 2024, 2023; Huang et al., [n. d.]). One of the most prominent concerns is the threat of model extraction (ME) attacks. As illustrated in Figure 1, such attacks construct a piracy model functionally equivalent or similar to the backend model of MLaaS, via querying the black-box APIs with carefully crafted inputs. Since its conceptualization in (Tramèr et al., 2016), a stream of work has been proposed to improve the efficacy of ME attacks (Tramèr et al., 2016; Orekondy et al., 2019a; Krishna et al., 2019; Chandrasekaran et al., 2020; Pal et al., 2019, 2020; Pengcheng et al., 2018; Shi et al., 2018; Juuti et al., 2019; Jagielski et al., 2020; Yu et al., 2020; Truong et al., 2021; Carlini et al., 2021; Papernot et al., 2016a, b, 2018, 2017).
@@
| Previous Conclusion | Refined Conclusion | Consistency |
| Identifying the architectures of victim models and using the same architectures in piracy models substantially boost the effectiveness of ME attacks (Chen et al., 2022b). | The architectures of piracy models have a limited impact on the performance of ME attacks, while more advanced models may not lead to more effective attacks. | |
| Using more complex models tends to achieve better attack performance (Shi et al., 2017; Krishna et al., 2019; Aïvodji et al., 2020). | The impact of model complexity varies with the concrete tasks (i.e., FER versus NLU) while pre-training is a more dominating factor. | |
| Semi-supervised learning substantially improves the query efficiency of ME attacks, especially when the query budget is low (Jagielski et al., 2020). | Semi-supervised learning improves the query efficiency of ME attacks, but the margin of improvement is not as large as reported in local experiments (Jagielski et al., 2020); further, it has a negative impact on adversarial fidelity. | |
| Active learning substantially improves the query efficiency of ME attacks (Pal et al., 2020). | Active learning yields only marginal improvements in query efficiency and can, in some cases, have a negative effect. | |
| Using adversarial examples improves both the query efficiency and the attack effectiveness (Papernot et al., 2017; Juuti et al., 2019; Pengcheng et al., 2018; Yu et al., 2020). | Using adversarial examples has a limited or even negative impact on attack fidelity, but may improve adversarial fidelity. | |
| Perturbing output confidence scores effectively mitigates ME attacks (Orekondy et al., 2019b; Lee et al., 2019). | Output quantization weakens ME attacks but is not sufficiently effective. |
– Leveraging MeBench, we conduct an empirical study on leading MLaaS platforms (i.e., Amazon, Microsoft, Face++, and Google) in the tasks of facial emotion recognition (FER) and natural language understanding (NLU). We show that today’s real-world MLaaS APIs still exhibit significant vulnerability to ME attacks, while the characteristics of vulnerability vary greatly across different platforms and tasks.
@@ 3. Status quo of ME vulnerability
Q1: How vulnerable are today’s MLaaS APIs to ME attacks?
Q2: How do various factors influence such vulnerability?
Q3: How does the vulnerability evolve over time?
Q4: How effective are the defenses (if any) employed on the MLaaS platforms?
@@ 3.1. Experiment setting
APIs – We consider the APIs of 4 leading MLaaS service providers: Amazon, Google, Microsoft, and Face++. By default, we assume the adversary has access to the complete query response (i.e., classification labels and confidence scores). Specifically, in the FER task, each query response contains the confidence scores of different expressions (e.g., “calm”). In the NLU task, each query response includes the confidence scores of different sentiments (e.g., “positive”). The details of each API are deferred to Table 14.
@@ 3.2.2. Comparison with prior work
MLaaS APIs – Among prior work, Cloudleak (Yu et al., 2020) conducts ME attacks against MLaaS APIs and measures attack fidelity with respect to different APIs and query budgets. Specifically, it reports 81% fidelity and 58% accuracy against the Face++ API on the KDEF dataset. Furthermore, it reports that the attacks against other platforms such as Microsoft and Clarifai, as well as different tasks including traffic sign recognition, flower, and NSFW, all attain over 82% fidelity and 70% accuracy. This clearly showcases the widespread nature of this vulnerability across various APIs and tasks.
In our experiments, on the same platform and dataset (i.e., KDEF and Face++), we observe that the attack attains similar results (74% fidelity and 66% accuracy), while the performance gap with (Yu et al., 2020) may be attributed to the scratch model and the difference of testing data. In (Yu et al., 2020), the testing data is synthesized, while we randomly sample 20% of the KDEF dataset as the testing set. Further, similar to (Yu et al., 2020), the ME vulnerability varies considerably across different APIs. In particular, the attack attains over 80% fidelity against the Microsoft API across all the datasets, while the other APIs demonstrate more variations from one dataset to another.
@@ 3.3.1. Optimizers
We specifically consider 4 popular optimizers: SGD (Robbins and Monro, 1951), Adam (Kingma and Ba, 2017), AdamW (Loshchilov and Hutter, 2019), and Lion (EvoLved Sign Momentum) (Chen et al., 2023a), which is a recently proposed optimizer. We configure the basic ME attack with varied optimizers and evaluate its effectiveness in the FER and NLU tasks.
@@
FER – We compare different optimizers using RAFDB on the Amazon and Microsoft APIs. Following (Chen et al., 2023a), we set the hyper-parameters (e.g., learning rate and decay rate ) for different optimizers as follows: SGD with and , Adam/AdamW with and , and Lion with and . We have the following key observations.
Table 3 shows the attack fidelity at the end of 200 training epochs. Under the same query budget, Lion, Adam, and AdamW attain comparable attack fidelity. There is no distinct advantage between Lion and AdamW; however, it is evident that SGD lags behind the performance of other optimizers significantly. We further examine the training dynamics of different optimizers. Figure 4 summarizes how the attack fidelity varies with the number of training epochs. Despite the claimed superior convergence speed of Lion in supervised learning (Chen et al., 2023a), we observe that Adam/AdamW actually converges much faster than Lion in the ME attack, while SGD converges fairly slowly.
NLU – We also set the hyper-parameters of optimizers following (Chen et al., 2023a): SGD with and , Adam/AdamW with and , and Lion with and .
Table 3 summarizes the attack performance corresponding to different optimizers on the IMDB dataset. Lion generally outperforms the other optimizers on both Microsoft and Amazon APIs. Our observations corroborate prior findings (Chen et al., 2023a) that Lion outperforms alternative optimizers in fine-tuning large language models.
@@ 3.3.2. Piracy models
Model family – For FER, we compare 5 popular architectures, GoogLeNet (Szegedy et al., 2014), DenseNet (Huang et al., 2018), ResNet (He et al., 2016), EfficientNet (Tan and Le, 2020), Alex -Net (Krizhevsky, 2014), using RAFDB on the Microsoft and Face++ APIs, with results shown in Figure 5. For NLU, we compare 2 popular LLMs, RoBERTa and XLNet, using IMDB on the Amazon and Microsoft APIs, with results summarized in Table 5. We have the following interesting findings.
@@
Model complexity – We further assess the impact of model complexity (within a single model family). In FER, we use ResNet as the piracy model and vary its width and depth, spanning across ResNet18, ResNet50, ResNet101, and Wide-ResNet50. As illustrated in Figure 6, the attack fidelity across different architectures differs by less than and on Amazon and Face++, respectively. Intuitively, as the model complexity reaches a certain level, further increasing the model width or depth does not lead to a notable improvement in attack performance. Similarly, in NLU, as shown in Table 5, for both RoBERTa and XLNet, using large models yields marginal improvement over base models. It is worth pointing out that compared with the base models, the large models of RoBERTa and XLNet are pre-trained using more data ( additional data) (Yang et al., 2019; Liu et al., 2019).
@@
In FER, we compare the attacks using a randomly initialized ResNet50 model as the piracy model with that using ResNet50 models pre-trained on public datasets. We consider two pre-training datasets, ImageNet-1K and VggFace2 (Cao et al., 2018). As shown in Figure 7, pre-training marginally improves the attack performance, while the improvement margin varies with the APIs. For instance, compared with VGGFace2, ImageNet-1K leads to more improvement on Microsoft but less so on Face++. In NLU, we compare randomly initialized XLNet and RoBERTa and that pre-trained on the BERT dataset (Devlin et al., 2018), with results shown in Table 6. Surprisingly, unlike the PER task, pre-training language models significantly improves the attack performance. For instance, the pre-training of RoBERTa boosts fidelity by over 25%.
@@ 3.3.3. Attack strategies
It is observed that against the Amazon API, using PGD or CW to enhance ME attacks leads to a considerable improvement in attack fidelity (around 5.61% on average) on FER+, while the effect on RAFDB is fairly mixed. Meanwhile, against Face++, adversarial learning-enhanced ME attack seems not only ineffective but even under-performs the basic attack. However, across all the cases, adversarial learning leads to a consistent improvement in adversarial fidelity. Our findings indicate that clean queries (sampled from the reference dataset) and adversarial queries (generated by adversarial attacks) seem to contribute to ME attacks differently, respectively improving the fidelity and adversarial fidelity of piracy models.
@@ 3.5. Q4: Defenses
It is unclear what specific defense mechanisms are employed by the MLaaS platforms. However, when examining their APIs, Amazon, Microsoft, and Face++ will provide exact confidence values for each class. In contrast, Google’s FER categorizes its output using five distinct likelihood descriptors, (i.e., “very_unlikely”, “unlikely”, “possible”, “likely” and “very_likely”) offering a different approach to presenting its analysis. This approach seems to be an attempt to obscure the confidence levels of the model’s output. The central concept behind such quantization is to reduce the granularity of the confidence values returned by the API, thereby making it more challenging for potential attackers to discern intricate details about the model’s internal processes. We mimicked Google’s approach and assess the effectiveness of quantization by applying it to other APIs. Specifically, We achieve this by discretizing the confidence values into intervals and substituting the confidence value with the midpoint of the corresponding interval (e.g., is represented as ). We manually reduce the granularity of the returned confidence information to examine its impact on the performance of the Model Extraction (ME) attack.
@@
| Face++ | +0.430.08 | -0.760.18 | +0.310.12 | -0.980.08 |
Table 13 presents the analysis results. When applied to the Amazon API, quantization resulted in an average decrease of 1.68% in the attack fidelity. In contrast, the Face++ API exhibits more mixed results in mitigating ME attacks.
@@ 6. Conclusion
generally improves fidelity but has a negative impact on adversarial fidelity, active learning provides a marginal enhancement, and surprisingly, adversarial learning may even negatively impact ME attacks. (v) The transferability of extracted piracy models depends more on the datasets, rather than other factors such as APIs, models, or ME attacks. (vi) API output quantization may weaken the effect of ME attacks, yet it remains inadequately effective.
@@ A.2. Datasets
| Task | Dataset | Size | Description |
| F E R | KDEF | 562*762 | The exaggerated facial expressions made by subjects |
| RAFDB | 100*100 | Great-diverse facial images downloaded from the Internet | |
| EXPW | 224*224 | Wild human facial expressions captured at social events | |
| FER+ | 48*48 | Black and white human face expressions | |
| N L U | IMDB | 234 words | Binary sentiment classification dataset, highly polar movie reviews collected from the internet movie database (IMDB) |
| Yelp | 140 words | Information from user reviews of different restaurants on Yelp. Ratings can be from 1-5, not a polarized dataset |